OpenStudy (anonymous):

how do you simplify i^17?

OpenStudy (shamil98):

when i is put to an odd exponent you always get -i. i^3 = -i i^5 = -i i^7 = -i and so on.

OpenStudy (anonymous):

so it's just -i?

OpenStudy (shamil98):

yep.

OpenStudy (solomonzelman):

NO

OpenStudy (solomonzelman):

\[i^{16}~\times i\]

OpenStudy (anonymous):

divide the exponent by 4 and look at the remainder. take i to that power and you'll have your answer

OpenStudy (shamil98):

oh oops i^17 = i forgot i^16 = 1

OpenStudy (solomonzelman):

it's i

OpenStudy (anonymous):

Ok, when you have even numbers you always get -1. When you have odd you always get i. Therefore, i^16 * i = -1 * i

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i^3 = -i i^4 = 1 i^2 = -1

OpenStudy (solomonzelman):

\[i^{4n}=1\]\[i^{4n+1}=i\]\[i^{4n+2}=-1\]\[i^{4n+3}=-i\]\[i^{4n+4}=i^{4n}=1\]YOU GET THE IDEA.

OpenStudy (anonymous):

ok thanks :)

OpenStudy (anonymous):

Snap! I messed up. Solomon is right.. srry

OpenStudy (solomonzelman):

Anytime!

OpenStudy (anonymous):

i^1 = i, i^2 = -1, i^3 = -i, i^4 = 1, etc The pattern repeats, grouping them into groups of four. Divide the exponent by 4 and take the remainder. Raise i to the power of remainder. Since 17 divide by 4 yields remainder of 1, i^17 = i^1 = i answer: i @kiergurl5